# Complete Characterizations of Well-Posedness in Parametric Composite Optimization

### v2

@@ 3 Qualification and Second-Order Variational Conditions
It follows from [Bonnans2000, Theorem 3.9 and Proposition 3.17] that the Lagrange multiplier set $\Lambda(x,0,0)$ is nonempty, convex, and compact if and only if RCQ holds at $x$. If $\bar{x}$ is a local minimizer of problem (1.1) and RCQ is satisfied at $\bar{x}$, then there exists a Lagrange multiplier $\bar{u}\in\mathcal{R}^{m}$ such that $(\bar{x},\bar{u})\in S_{\operatorname{KKT}}(0,0)$. It is said that the strong Robinson constraint qualification (SRCQ) is satisfied at $\bar{x}$ with respect to $\bar{u}$ if we have
@@
by using the generalized Jacobian of the proximal mapping (2.3). Following [Tangwang2024], consider the generalized version of the strong second-order sufficient condition for the composite optimization problem (1.1) that extends the classical one for NLPs introduced in [Robinson1980].
@@
It is shown in [Tangwang2024, Proposition 3.1] that every $\mathcal{C}^{2}$-cone reducible function satisfies assumption (A1). Furthermore, it is proved in Appendices of [MordukhovichTangWang2025, MordukhovichTangWang-tilt-2025] that assumption (A2) holds for the indicator function of the positive semidefinite cone, the nuclear norm function, and the spectral norm function, which all are important in applications.
@@
Since $d^{2}f(\bar{x},\bar{\mu})$ is positive homogeneous of degree 2, we only need to prove the result when $\|d\|\leq 1$. It follows from [Rockafellar1998, Proposition 13.20 and Theorem 13.40] and [MordukhovichTangWang-tilt-2025, Proposition 3.2] that $d^{2}f(x^{k},\mu^{k})$ is convex and $D(\partial f)(x^{k},\mu^{k})(d^{k})=\partial\left(\frac{1}{2}d^{2}f(x^{k},\mu^{k})\right)(d^{k})$ is bounded. Picking $v^{k}\in D(\partial f)(x^{k},\mu^{k})(d^{k})$, we get by (2.11) and (2.13) that $d^{k}\in D\operatorname{Prox}_{f}(x^{k}+\mu^{k})(v^{k}+d^{k})=\{\operatorname{Prox}_{\frac{1}{2}d^{2}f(x^{k},\mu^{k})}(v^{k}+d^{k})\}$. Then $\operatorname{Prox}_{f}$ is directionally differentiable at $x^{k}+\mu^{k}$ in the direction $v^{k}+d^{k}$ and there exists $\tau_{t}\downarrow 0$ as $t\rightarrow\infty$ such that
@@
Letting $k_{t_{j}}\rightarrow\infty$ above, it follows from the directional differentiability of $\operatorname{Prox}_{f}$ at $\bar{x}+\bar{\mu}$ in the direction $\bar{v}+d$ that $D\operatorname{Prox}_{f}(\bar{x}+\bar{\mu})(\bar{v}+d)=\{d\}$. Combining (2.11) and (2.13), we get $\bar{v}\in D(\partial f)(\bar{x},\bar{\mu})(d)$. Applying finally [Tangwang2024, Proposition 3.4] yields
@@
The next lemma reveals a fundamental relationship between the second subderivative (2.5) of an extended-real-valued convex function and its second-order subdifferential introduced in [Mordukhovich1992] as the limiting coderivative in (2.9) of the limiting subdifferential (2.2).
@@
It follows from assumption (A1) and [MordukhovichTangWang2025, Lemma 5.2] that
@@
Applying [Mohammadi2020, Proposition 3.8], [Chieu2021, Lemma 3.6], and [Rockafellar1998, Theorem 13.57] tells us that $f$ is twice epi-differentiable at $\bar{x}$ for $\bar{\mu}$ and the inequality $\leq$” holds in (3.11). To verify the reverse inequality therein, deduce from [MordukhovichTangWang-tilt-2025, Theorem 3.1] the existence of sequences $(x^{k},\mu^{k})\rightarrow(\bar{x},\bar{\mu})$ with $\mu^{k}\in\partial f(x^{k})$ and $d^{k}\rightarrow d$ as $k\to\infty$ along which we have
@@
It follows from [Mohammadi2020, Theorem 3.8] and [MordukhovichTangWang-tilt-2025, Proposition 3.2] that
@@
Furthermore, [MordukhovichTangWang-tilt-2025, Proposition 3.3] tells that
@@
Applying the result in [MordukhovichTangWang-tilt-2025, Appendix], SSOSC (3.10) can be expressed as the following form according to the value of the nuclear norm $\|A\|_{*}$, which is the sum of the singular values of $A$.
@@ 4 Characterizations of Second-Order Qualification Condition
The second-order qualification condition (SOQC) in problems of composite optimization of type (1.1) at $\bar{x}$ for $\bar{u}$ was introduced in [Mordukhovich2012] in the form
@@
where $(\bar{x},\bar{u})$ is a solution to the perturbed KKT system (1.3) with $(a,b)=(0,0)$. As shown in [Mordukhovich2023, Lemma 7.2], SOQC (4.1) reduces to the classical linear independence constraint qualification in the case of smooth problems in nonlinear programming. The main result of this section provides complete characterizations of SOQC for (1.1) when the function $g$ is ${\cal C}^{2}$-cone reducible.
@@
$f:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}$ be an l.s.c. convex function with $\bar{x}\in\operatorname{dom}\,f$ and $\bar{\mu}\in\partial f(x)$ under the fulfillment of assumption (A1). Then we have the representation
$D^{*}(\partial f)(\bar{x},\bar{\mu})(0)=\big\{d\,\big|\,(d,0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(\bar{\mu},-1))(0,0)\big\}.$ (4.2)
@@
It follows from [Rockafellar1998, Theorem 8.9] that for any $x\in\operatorname{dom}f$, $\mu\in\partial f(x)$, and $\lambda>0$, we get $\lambda(\mu,-1)\in\partial\delta_{\operatorname{epi}f}(x,f(x))$. Using assumption (A1), [Mohammadi2020, Theorem 3.8], [Rockafellar1998, Example 13.62], and Proposition 2.1 above tell us that $\delta_{\operatorname{epi}f}$ is parabolically regular at $(x,f(x))$ for $\lambda(\mu,-1)$ and parabolically epi-differentiable at $(x,f(x))$ for every $(z,\gamma)\in\operatorname{dom}d\delta_{\operatorname{epi}f}(x,f(x))$. Then it follows from [MordukhovichTangWang2025, Lemma 5.2] and Proposition 2.1 that we have the equalities
@@
Pick any $(\hat{y},0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x}),(\bar{\mu},-1)))(0,0)$ and deduce from the definitions in (2.9) and (2.6) the existence of $((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))\in\operatorname{gph}\mathcal{N}_{\operatorname{epi}f}$ with $((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))\rightarrow((\bar{x},f(\bar{x})),(\bar{\mu},-1))$, $(d^{k},\beta^{k})\rightarrow(0,0)$, and $(y^{k},\gamma^{k})\in\widehat{D}^{*}\mathcal{N}_{\operatorname{epi}f}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))(d^{k},\beta^{k})$ such that $(y^{k},\gamma^{k})\rightarrow(\hat{y},0)$. Applying further [Rockafellar1998, Proposition 8.37] to any pair $(y^{k},\gamma^{k})\in\widehat{D}^{*}\mathcal{N}_{\operatorname{epi}f}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))(d^{k},\beta^{k})$ leads us to the estimate
@@
$\mbox{for all }\ ((\widetilde{y}^{k},\widetilde{\gamma}^{k}),(\widetilde{d}^{k},\widetilde{\beta}^{k}))\in\mathcal{T}^{k}:=\mathcal{T}_{\operatorname{gph}\mathcal{N}_{\operatorname{epi}f}}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1)),$
@@
It follows from (4.6) that $\widetilde{y}^{k}\in\mathcal{K}(x^{k},\mu^{k})\cap\mathbb{B}_{\mathcal{R}^{m}}$, and therefore
$\big\{d\,\big|\,(d,0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(\bar{\mu},-1))(0,0)\big\}$
@@
$\left.((\widetilde{y}^{k},\widetilde{\gamma}^{k}),(\widetilde{d}^{k},\widetilde{\beta}^{k}))\in\mathcal{T}^{k}\cap\mathbb{B}_{\mathcal{R}^{m+1}\times\mathcal{R}^{m+1}},\,\gamma^{k}\rightarrow 0,\,(d^{k},\beta^{k})\rightarrow(0,0)\right\}$
@@
where $\mathop{{\rm Lim}\,{\rm sup}}$” signifies the limits of convergent subsequences (Painvlevé-Kuratowski outer limit); cf. (2.6) and (2.9). On the other hand, we get for $D^{*}(\partial f)(\bar{x},\bar{\mu})(0)$ by using (4.3) that
[…]
This verifies (4.2) and thus completes the proof of the lemma. ∎
@@
Applying [Rockafellar1970, Corollary 4.7.2] tells us that
@@
Based on the above lemmas, we now establish relationships between the second-order qualification condition, constraint nondegeneracy, and Aubin property of the multiplier mapping.
@@
Let $(\bar{x},\bar{u})$ solve the KKT system (1.3) with $(a,b)=(0,0)$. Denote $M_{\bar{x}}(a,b):=\Lambda(\bar{x},a,b)$ for the Lagrange multiplier mapping (3.2) and consider the following assertions:
(i) The Lagrange multiplier mapping $M_{\bar{x}}$ is Lipschitz-like at the origin for $\bar{u}$.
(ii) SOQC (4.1) is satisfied at $\bar{x}$ for $\bar{u}$.
(iii) The constraint nondegeneracy condition (3.4) holds at $\bar{x}$ for $\bar{u}$.
Then we have the following relationships:
[…]
@@
First we compute the coderivative of $M_{\bar{x}}$ at $(0,0)$ for $\bar{u}$. To proceed, define $\Phi_{\bar{x}}(u):=\Phi(\bar{x},u)$ and observe that $M_{\bar{x}}=\Phi_{\bar{x}}^{-1}$, which means that $(\Delta\xi,\Delta\eta)\in D^{*}\Phi_{\bar{x}}(\bar{u},(0,0))(-d)$ for any $d\in D^{*}M_{\bar{x}}((0,0),\bar{u})(\Delta\xi,\Delta\eta)$. It follows from [Mordukhovich2006, Theorem 1.62] that
$D^{*}\Phi_{\bar{x}}(\bar{u},(0,0))(-d)=\left[\begin{array}[]{c}-\nabla F(\bar{x})^{T}d\\ D^{*}(\partial g^{*})(\bar{u},F(\bar{x}))(-d)\end{array}\right].$
Consequently, we have that $d\in D^{*}M_{\bar{x}}((0,0),\bar{u})(0,0)$ if and only if $d$ satisfies the conditions
$-\nabla F(\bar{x})^{T}d=0\;\mbox{ and }\;0\in D^{*}(\partial g^{*})(\bar{u},F(\bar{x}))(-d).$
[…]
@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
It follows from [Rockafellar1998, Theorem 8.9] that $\widetilde{S}_{\operatorname{KKT}}((0,0),(0,0))=\{((x,g(F(x))),(u,-1))\,|\,(x,u)\in S_{\operatorname{KKT}}(a,0)\}$, and hence $\widetilde{S}_{\operatorname{KKT}}((0,0),(0,0))$ can be seen as the set of solutions to the KKT system associated with the canonically perturbed constrained optimization problem given by
@@
Proposition 2.1, [MordukhovichTangWang2025, Lemma 5.2], [MordukhovichTangWang-tilt-2025, Proposition 3.2], and [Rockafellar1998, Theorem 13.40] ensure that
@@
Combining Proposition 2.1, [Bonnans2000, Proposition 3.41], the parabolic regularity of $\delta_{\operatorname{epi}g}$ and $g$ yields
@@
It follows from Lemma 2.2 that the solution mapping $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if the modified KKT mapping $\widetilde{S}_{\operatorname{KKT}}$ has this property at $((0,0),(0,0))$ for ((x,c),(u,1))((x,c),(u,-1)). By the above expression of $d^{2}\delta_{\operatorname{epi}g}$, the original SOSC (3.6) for (5.2) is equivalent to the SOSC version in (5.1). The equivalence between SRCQ for (5.2) and the original SRCQ (3.3) is trivial. Due to the fact that $(x,g(F(x)))$ is also the unique local minimizer of problem (5.2), the proof is completed by applying [MohammadiMordukhovichSarabi2020, Theorem 4.1]. ∎
Recall yet another Lipschitzian stability property for local minimizers of extended-real-valued functions $\varphi\colon\mathcal{R}^{n}\rightarrow\overline{\mathcal{R}}$. Following [Poliquin1998], we say that $\bar{x}\in{\rm dom}\,\varphi$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$ if there exists a number $\gamma>0$ such that the mapping
@@
is single-valued and Lipschitz continuous with modulus $\kappa$ on some neighborhood of the origin with $M_{\gamma}(0)=\{\bar{x}\}$. Over the years, this property has been largely studied and applied in variational analysis and optimization; see, e.g., the recent book [Mordukhovich2024] for major achievements in this direction. The following observation tells us that the tilt stability of a local minimizer $\bar{x}$ for a parabolically regular function $\varphi$ yields the second-order subdifferential condition (5.3) when the corresponding multiplier sets around $\bar{x}$ are singletons.
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ being a corresponding Lagrange multiplier. Suppose that RCQ (3.1) holds at $\bar{x}$, that $g$ satisfies (A1) at $F(\bar{x})$, and that there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that $\Lambda(x,a,0)$ is a singleton for $x\in\mathcal{U}$, $a\in\mathcal{V}$, and $(x,u)\in S_{\operatorname{KKT}}(a,0)$. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$, then we have
@@
Take $\bar{z}\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)$ with $\min\limits_{z\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)}\left\langle\nabla F(\bar{x})d,z\right\rangle=\left\langle\bar{z},\nabla F(\bar{x})d\right\rangle$ and find sequences $(x^{k},u^{k})\rightarrow(\bar{x},\bar{u})$ with $u^{k}\in\partial g(F(x^{k}))$ and $(z^{k},d^{k})\rightarrow(\bar{z},d)$ with $z^{k}\in\widehat{D}^{*}(\partial g)(F(x^{k}),u^{k})(\nabla F(x^{k})d^{k})$ satisfying $\nabla h(x^{k})+\nabla F(x^{k})^{T}u^{k}\in\mathcal{V}$ and $\Lambda(x^{k},\nabla h(x^{k})+\nabla F(x^{k})^{T}u^{k},0)=\{u^{k}\}$ when $k$ is sufficiently large. It follows from [MordukhovichTangWang-tilt-2025, Lemma 3.1], (2.11), (2.13), and [MordukhovichTangWang2025, Lemma 5.2] that $\nabla F(x^{k})d^{k}\in\operatorname{dom}D(\partial g)(F(x^{k}),u^{k})$. Combining [MordukhovichTangWang-tilt-2025, Lemma 4.1] and [Chieu2018, Theorem 3.3] brings us to the inequality
@@
Using [Rockafellar1998, Proposition 8.37] and [Chieu2021, Lemma 3.6] implies that
@@
Using the above constructions, we recall the fundamental properties of (Lipschitzian) full stability. The first part of the following definition concerns primal full stability for local minimizers of the original problem (1.1) under perturbations in (1.2) introduced in [LevyPoliquinRockafellar2000], while the second notion of primal-dual full stability, which has been recently introduced in [BenkoRockafellar2024], addresses behavior of solution-multiplier pairs satisfying the associated KKT system (1.3) under perturbations.
@@
We refer the reader to [Mordukhovich2024, Chapter 5] for the state-of-the-art in the study of full stability of local minimizers from Definition 5.1(i) and to [Mordukhovich2024, Chapter 6] for full stability of parametric variational systems related to Definition 5.1(ii).

### v3

@@ 3 Qualification and Second-Order Variational Conditions
It follows from [4, Theorem 3.9 and Proposition 3.17] that the Lagrange multiplier set $\Lambda(x,0,0)$ is nonempty, convex, and compact if and only if RCQ holds at $x$. If $\bar{x}$ is a local minimizer of problem (1.1) and RCQ is satisfied at $\bar{x}$, then there exists a Lagrange multiplier $\bar{u}\in\mathcal{R}^{m}$ such that $(\bar{x},\bar{u})\in S_{\operatorname{KKT}}(0,0)$. It is said that the strict Robinson constraint qualification (SRCQ) is satisfied at $\bar{x}$ with respect to $\bar{u}$ if we have
@@
by using the generalized Jacobian of the proximal mapping (2.3). Following [42], consider the generalized version of the strong second-order sufficient condition for the composite optimization problem (1.1) that extends the classical one for NLPs introduced in [38].
@@
It is shown in [42, Proposition 3.1] that every $\mathcal{C}^{2}$-cone reducible function satisfies assumption (A1). Furthermore, it is proved in Appendices of [28, 29] that assumption (A2) holds for the indicator function of the positive semidefinite cone, the nuclear norm function, and the spectral norm function, which all are important in applications.
@@
Since $d^{2}f(\bar{x},\bar{\mu})$ is positive homogeneous of degree 2, we only need to prove the result when $\|d\|\leq 1$. It follows from [39, Proposition 13.20 and Theorem 13.40], [29, Proposition 3.2] and [28, Lemma 5.2] that $d^{2}f(x^{k},\mu^{k})$ is convex and $D(\partial f)(x^{k},\mu^{k})(d^{k})=\partial\left(\frac{1}{2}d^{2}f(x^{k},\mu^{k})\right)(d^{k})$ is bounded. Picking $v^{k}\in D(\partial f)(x^{k},\mu^{k})(d^{k})$, we get by (2.11) and (2.13) that $d^{k}\in D\operatorname{Prox}_{f}(x^{k}+\mu^{k})(v^{k}+d^{k})=\{\operatorname{Prox}_{\frac{1}{2}d^{2}f(x^{k},\mu^{k})}(v^{k}+d^{k})\}$. Then $\operatorname{Prox}_{f}$ is directionally differentiable at $x^{k}+\mu^{k}$ in the direction $v^{k}+d^{k}$ and there exists $\tau_{t}\downarrow 0$ as $t\rightarrow\infty$ such that
@@
Letting $k_{t_{j}}\rightarrow\infty$ above, it follows from the directional differentiability of $\operatorname{Prox}_{f}$ at $\bar{x}+\bar{\mu}$ in the direction $\bar{v}+d$ that $D\operatorname{Prox}_{f}(\bar{x}+\bar{\mu})(\bar{v}+d)=\{d\}$. Combining (2.11) and (2.13), we get $\bar{v}\in D(\partial f)(\bar{x},\bar{\mu})(d)$. Applying finally [42, Proposition 3.4] yields
@@
The next lemma reveals a fundamental relationship between the second subderivative (2.5) of an extended-real-valued convex function and its second-order subdifferential introduced in [30] as the limiting coderivative in (2.9) of the limiting subdifferential (2.2).
@@
It follows from assumption (A1) and [28, Lemma 5.2] that
@@
Applying [23, Proposition 3.8], [8, Lemma 3.6], and [39, Theorem 13.57] tells us that $f$ is twice epi-differentiable at $\bar{x}$ for $\bar{\mu}$ and the inequality $\leq$” holds in (3.11). To verify the reverse inequality therein, deduce from [29, Theorem 3.1] the existence of sequences $(x^{k},\mu^{k})\rightarrow(\bar{x},\bar{\mu})$ with $\mu^{k}\in\partial f(x^{k})$ and $d^{k}\rightarrow d$ as $k\to\infty$ along which we have
@@
It follows from [23, Theorem 3.8] and [29, Proposition 3.2] that
@@
Furthermore, [29, Proposition 3.3] tells that
@@
Applying the result in [29, Appendix], SSOSC (3.10) can be expressed as the following form according to the value of the nuclear norm $\|A\|_{*}$, which is the sum of the singular values of $A$.
@@ 4 Characterizations of Second-Order Qualification Condition
The second-order qualification condition (SOQC) in problems of composite optimization of type (1.1) at $\bar{x}$ for $\bar{u}$ was introduced in [25] in the form
@@
where $(\bar{x},\bar{u})$ is a solution to the perturbed KKT system (1.3) with $(a,b)=(0,0)$. As shown in [15, Lemma 7.2], SOQC (4.1) reduces to the classical linear independence constraint qualification in the case of smooth problems in nonlinear programming. The main result of this section provides complete characterizations of SOQC for (1.1) when the function $g$ is ${\cal C}^{2}$-cone reducible.
@@
$f:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}$ be an l.s.c. convex function with $\bar{x}\in\operatorname{dom}\,f$ and $\bar{\mu}\in\partial f(\bar{x})$ under the fulfillment of assumption (A1). Then we have the representation
$D^{*}(\partial f)(\bar{x},\bar{\mu})(d)$ $=\big\{y\,\big|\,(y,0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(\bar{\mu},-1))(d,\beta)\big\}.$ (4.2)
@@
It follows from [39, Theorem 8.9] that for any $x\in\operatorname{dom}f$, $\mu\in\partial f(x)$, and $\lambda>0$, we get $\lambda(\mu,-1)\in\partial\delta_{\operatorname{epi}f}(x,f(x))$. Using assumption (A1), [23, Theorem 3.8], [39, Example 13.62], and Proposition 2.1 above tell us that $\delta_{\operatorname{epi}f}$ is parabolically regular at $(x,f(x))$ for $\lambda(\mu,-1)$ and parabolically epi-differentiable at $(x,f(x))$ for every $(z,\gamma)\in\operatorname{dom}d\delta_{\operatorname{epi}f}(x,f(x))$. Then it follows from Proposition 2.1, [28, Lemma 5.2], [29, Proposition 3.2], and [39, Theorem 13.40] that we have the equalities
$D\mathcal{N}_{\operatorname{epi}f}((x,f(x)),(\mu,-1))(y,\gamma)=\frac{1}{2}\partial d^{2}\delta_{\operatorname{epi}f}((x,f(x)),(\mu,-1))(y,\gamma),$
$D(\partial f)(x,\mu)(y)=\frac{1}{2}\partial d^{2}f(x,\mu)(y),$
@@
Combining Proposition 2.1, [4, Proposition 3.41], the parabolic regularity of $\delta_{\operatorname{epi}f}$ and $f$ yields
[…]
Pick any $(y,0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(\bar{\mu},-1))(d,\beta)$ and deduce from the definitions in (2.9) and (2.6) the existence of $((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))\in\operatorname{gph}\mathcal{N}_{\operatorname{epi}f}$ with $((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))\rightarrow((\bar{x},f(\bar{x})),(\bar{\mu},-1))$, $(d^{k},\beta^{k})\rightarrow(d,\beta)$, and $(y^{k},\gamma^{k})\in\widehat{D}^{*}\mathcal{N}_{\operatorname{epi}f}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))(d^{k},\beta^{k})$ such that $(y^{k},\gamma^{k})\rightarrow(y,0)$. Applying further [39, Proposition 8.37] to any pair $(y^{k},\gamma^{k})\in\widehat{D}^{*}\mathcal{N}_{\operatorname{epi}f}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1))(d^{k},\beta^{k})$ leads us to the estimate
@@
$\qquad\mbox{for all }\ ((\widetilde{y}^{k},\widetilde{\gamma}^{k}),(\widetilde{d}^{k},\widetilde{\beta}^{k}))\in\mathcal{T}^{k}:=\mathcal{T}_{\operatorname{gph}\mathcal{N}_{\operatorname{epi}f}}((x^{k},f(x^{k})),\lambda^{k}(\mu^{k},-1)),$
@@
It follows from (4.6) that $\widetilde{y}^{k}\in\mathcal{K}(x^{k},\mu^{k})\cap\mathbb{B}_{\mathcal{R}^{m}}$, and therefore by (4.3) and (4.7)
$\big\{y\,\big|\,y,0)\in D^{*}\mathcal{N}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(\bar{\mu},-1))(d,\beta)\big\}$
@@
$\qquad\left.((\widetilde{y}^{k},\widetilde{\gamma}^{k}),(\widetilde{d}^{k},\widetilde{\beta}^{k}))\in\mathcal{T}^{k}\cap\mathbb{B}_{\mathcal{R}^{m+1}\times\mathcal{R}^{m+1}},\,\gamma^{k}\rightarrow 0,\,(d^{k},\beta^{k})\rightarrow(d,\beta)\right\}$
@@
$=\mathop{{\rm Lim}\,{\rm sup}}_{k\rightarrow\infty}\widehat{D}^{*}(\partial f)(x^{k},\mu^{k})(d^{k})$
$=D^{*}(\partial f)(\bar{x},\bar{\mu})(d),$
where $\mathop{{\rm Lim}\,{\rm sup}}$” signifies the limits of convergent subsequences (Painlevé-Kuratowski outer limit); cf. (2.6) and (2.9). This verifies (4.2) and thus completes the proof of the lemma. ∎
@@
Applying [40, Corollary 4.7.2] tells us that
@@
Based on the above lemmas, we now establish relationships between the second-order qualification condition and constraint nondegeneracy.
@@
Let $(\bar{x},\bar{u})$ solve the KKT system (1.3) with $(a,b)=(0,0)$. If the constraint nondegeneracy condition (3.4) holds at $\bar{x}$ for $\bar{u}$ and in addition $g$ is $\mathcal{C}^{2}$-cone reducible, then SOQC (4.1) is satisfied at $\bar{x}$ for $\bar{u}$.
@@
This follows from Lemma 4.2 combined with [26, Corollary 3.8]. ∎
@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
It follows from [39, Theorem 8.9] that $\widetilde{S}_{\operatorname{KKT}}((0,0),(0,0))=\{((x,g(F(x))),(u,-1))\,|\,(x,u)\in S_{\operatorname{KKT}}(a,0)\}$, and hence $\widetilde{S}_{\operatorname{KKT}}((0,0),(0,0))$ can be seen as the set of solutions to the KKT system associated with the canonically perturbed constrained optimization problem given by
@@
Proposition 2.1, [28, Lemma 5.2], [29, Proposition 3.2], and [39, Theorem 13.40] ensure that
@@
Combining Proposition 2.1, [4, Proposition 3.41], the parabolic regularity of $\delta_{\operatorname{epi}g}$ and $g$ yields
@@
It follows from Lemma 2.2 that the solution mapping $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(a,0)$ for $(x,u)$ if and only if the modified KKT mapping $\widetilde{S}_{\operatorname{KKT}}$ has this property at $((0,0),(0,0))$ for ((x,c),(u,1))((x,c),(u,-1)). By the above expression of $d^{2}\delta_{\operatorname{epi}g}$, the original SOSC (3.6) for (5.2) is equivalent to the SOSC version in (5.1). The equivalence between SRCQ for (5.2) and the original SRCQ (3.3) is trivial. Due to the fact that $(x,g(F(x)))$ is also the unique local minimizer of problem (5.2), the proof is completed by applying [20, Theorem 4.1]. ∎
Recall yet another Lipschitzian stability property for local minimizers of extended-real-valued functions $\varphi\colon\mathcal{R}^{n}\rightarrow\overline{\mathcal{R}}$ in (1.1). Following [37], we say that $\bar{x}\in{\rm dom}\,\varphi$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$ if there exists a number $\gamma>0$ such that the mapping
@@
is single-valued and Lipschitz continuous with modulus $\kappa$ on some neighborhood of the origin with $M_{\gamma}(0)=\{\bar{x}\}$. Over the years, this property has been largely studied and applied in variational analysis and optimization; see, e.g., the recent book [34] for major achievements in this direction. The following observation tells us that the tilt stability of a local minimizer $\bar{x}$ for a parabolically regular function $\varphi$ yields the second-order subdifferential condition (5.3) when the corresponding multiplier sets around $\bar{x}$ are singletons.
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ being a corresponding Lagrange multiplier. Suppose that RCQ (3.1) holds at $\bar{x}$, that $g$ satisfies (A1) at $F(\bar{x})$, and that there exist neighborhoods $\mathcal{U}$ of $\bar{x}$ and $\mathcal{V}$ of the origin such that $\Lambda(x,a,0)$ is a singleton for $x\in\mathcal{U}$, $a\in\mathcal{V}$, and $(x,u)\in S_{\operatorname{KKT}}(a,0)$. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa>0$, then we have the second-order condition
@@
Take $\bar{z}\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)$ with $\min\limits_{z\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)}\left\langle\nabla F(\bar{x})d,z\right\rangle=\left\langle\bar{z},\nabla F(\bar{x})d\right\rangle$ and find sequences $(x^{k},u^{k})\rightarrow(\bar{x},\bar{u})$ with $u^{k}\in\partial g(F(x^{k}))$ and $(z^{k},d^{k})\rightarrow(\bar{z},d)$ with $z^{k}\in\widehat{D}^{*}(\partial g)(F(x^{k}),u^{k})(\nabla F(x^{k})d^{k})$ satisfying $\nabla h(x^{k})+\nabla F(x^{k})^{T}u^{k}\in\mathcal{V}$ and $\Lambda(x^{k},\nabla h(x^{k})+\nabla F(x^{k})^{T}u^{k},0)=\{u^{k}\}$ when $k$ is sufficiently large. It follows from [29, Lemma 3.1], (2.11), (2.13), and [28, Lemma 5.2] that $\nabla F(x^{k})d^{k}\in\operatorname{dom}D(\partial g)(F(x^{k}),u^{k})$. Combining [29, Lemma 4.1] and [9, Theorem 3.3] brings us to the inequality
@@
Using [39, Proposition 8.37] and [8, Lemma 3.6] implies that
@@
Using the above constructions, we recall the fundamental properties of (Lipschitzian) full stability. The first part of the following definition concerns primal full stability for local minimizers of the original problem (1.1) under perturbations in (1.2) introduced in [18], while the second notion of primal-dual full stability, which has been recently introduced in [2], addresses behavior of solution-multiplier pairs satisfying the associated KKT system (1.3) under perturbations.
@@
We refer the reader to [34, Chapter 5] for the state-of-the-art in the study of full stability of local minimizers from Definition 5.1(i) and to [34, Chapter 6] for full stability of parametric variational systems related to Definition 5.1(ii).